A body falls freely through a distance of  10 m from rest. It is then brought to rest in 1 second by a vertical force. A similar second force can bring the body to rest in 2 seconds. Show that the first force is 17/12 times the second force. 


1
Expert's answer
2021-05-08T14:39:11-0400
"mv=Ft\\F_1t_1=F_2t_1\\\frac{F_1}{F_2}=\frac{t_2}{t_1}=\frac{2}{1}=2"